# A Comprehensive Survey and Taxonomy on Point Cloud Registration Based on Deep Learning

### v1

@@ 1 Introduction
With the progress of sensor technology, acquiring high-precision point cloud data has become more accessible and prevalent (Zhang et al., 2023b; Uy et al., 2019). Point cloud registration (PCR), as a pivotal tool in point cloud data processing, aims to align the point cloud data with a common coordinate system, enabling precise three-dimensional (3D) modeling (Qin et al., 2023; Chen et al., 2022b). This registration process establishes a dependable foundation for point cloud analysis and various applications (Huang et al., 2021b).
@@
The differences between our survey and others. (Gu et al., 2020) only review traditional PCR algorithms, DL-based algorithms are not involved. (Zhang et al., 2020) and (Huang et al., 2021b) conduct a summary of DL-based PCR algorithms. However, recent advances in unsupervised algorithms are not elaborated. Additionally, they did not provide a comprehensive overview of the latest research developments in the PCR field. To address these gaps, our study conducts a comprehensive survey and taxonomy of DL-based supervised and unsupervised registration algorithms. The taxonomy is concisely summarized in Figure 1, providing a clear and structured overview of the PCR algorithms.
@@ 2.1 Definition
PCR Supervised Descriptor Extraction Two-view Multi-view Keypoint-based Keypoint-free Correspondence Search Full-object Partial-object Overlap Prediction Optimizing the Similarity Matrix Outlier Filtering Transformation Parameter Estimation Optimization ICP-based Probabilistic -based Multimodal DeepVCP (Lu et al., 2019a), 3DSmoothNet (Gojcic et al., 2019), Deng et al. (Deng et al., 2019), HRegNet (Lu et al., 2021), SpinNet (Ao et al., 2021), StickyPillars (Fischer et al., 2021), YOHO (Wang et al., 2022a), RoReg (Wang et al., 2023b), BUFFER (Ao et al., 2023). OIF-PCR (Yang et al., 2022), GeDi (Poiesi and Boscaini, 2022), GeoTransformer (Qin et al., 2023), RoITr (Yu et al., 2023a). MVDesc (Zhou et al., 2018), Li et al. (Li et al., 2020b), Gojcic et al. (Gojcic et al., 2020), Wang et al. (Wang et al., 2023a). PointNetLK (Aoki et al., 2019), DCP (Wang and Solomon, 2019a), PointNetLK Revisited (Li et al., 2021). Predator (Huang et al., 2021a), OMNet (Xu et al., 2021), PCAM (Cao et al., 2021), STORM (Wang et al., 2022b), REGTR (Yew and Lee, 2022). PRNet (Wang and Solomon, 2019b), RPMNet (Yew and Lee, 2020), FIRE-Net (Wu et al., 2021) 3DRegNet (Pais et al., 2020), DHVR (Lee et al., 2021), PointDSC (Bai et al., 2021), DLF (Zhang et al., 2022a), Chen et al.(Chen et al., 2022b), MAC (Zhang et al., 2023a). DetarNet (Chen et al., 2022c), FINet (Xu et al., 2022). DCP (Wang and Solomon, 2019a), PRNet (Wang and Solomon, 2019b), IDAM (Li et al., 2020a). DeepGMR (Yuan et al., 2020), OGMM (Mei et al., 2023a), Chen et al.(Chen et al., 2023a), VBReg (Jiang et al., 2023). PCR-CG (Zhang et al., 2022c), ImLoveNet (Chen et al., 2022a), IMFNET (Huang et al., 2022c), GMF (Huang et al., 2022b), PEAL (Yu et al., 2023b). Unsupervised Correspondence -free PPF-FoldNet (Deng et al., 2018), PCRNet (Sarode et al., 2019), UPCR (Zhang et al., 2021), UGMM (Huang et al., 2022a), Sun et al. (Sun et al., 2023). Correspondence -based CEMNet (Jiang et al., 2021), RIENet (Shen et al., 2022), UDPReg (Mei et al., 2023b).
Figure 1: A taxonomy of PCR algorithms.
@@ 2.2 Datasets
Datasets for PCR can be broadly classified into two categories: artificially synthesized and acquired through real instruments. Each category exhibits unique characteristics and has different levels of applicability in PCR tasks. Synthesized point cloud datasets are typically composed of virtual models created by using computer graphics techniques to replicate real-world environments. These datasets contain the object-level ModelNet40 (Wu et al., 2015) and ShapeNet (Chang et al., 2015), which consist of data generated through computer-aided design, as well as the scene-level dataset ICL-NUIM (Choi et al., 2015) and FlyingShapes (Chen et al., 2023b).
However, while synthetic data are beneficial for certain applications, they often lack the complexity and variability found in real-world scenarios. Consequently, incorporating real-world data into training and validation processes is essential for obtaining robust algorithms. Datasets comprising realistic point cloud data include Stanford (Curless and Levoy, 1996), ETH (Pomerleau et al., 2012), KITTI (Geiger et al., 2012), Apollo-SouthBay (Lu et al., 2019b), ScanObjectNN (Uy et al., 2019), and WHU-TLS (Dong et al., 2020). Furthermore, there is a 3DMatch (Zeng et al., 2017) dataset that comprises both synthesized and realistic scans. The attributes of various datasets are summarized in Table 1.
@@ 2.3 Metrics
Stanford (Curless and Levoy, 1996) Real 10 O
ETH (Pomerleau et al., 2012) Real 36 S
KITTI (Geiger et al., 2012) Real 22 S
ModelNet40 (Wu et al., 2015) Syn 12311 O
ShapeNet (Chang et al., 2015) Syn 55000+ O
ICL-NUIM (Choi et al., 2015) Syn 8 S
3DMatch (Zeng et al., 2017) Syn&Real 62 S
Apollo-SouthBay (Lu et al., 2019b) Real 6 S
ScanObjectNN (Uy et al., 2019) Real 15000+ O
WHU-TLS (Dong et al., 2020) Real 115 S
FlyingShapes (Chen et al., 2023b) Syn 200 S
Table 1: Datasets for PCR tasks. Syn means synthetic point clouds. Real means realistic point clouds. S and O denote scene-level and object-level, respectively.
@@ 3.1 Descriptor Extraction
3.1.1 Two-view
The first perspective involves two-view registration, which emerges as the prevalent approach in the field of PCR. We further classify these methods into two categories: keypoint-based and keypoint-free.
@@ 3.1.1 Two-view
Firstly, points, serving as the fundamental elements of point clouds, are discrete and unlinked 3D entities. Consequently, extracting descriptors from points typically necessitates the construction of intricate local relationships. In DeepVCP (Lu et al., 2019a), point weighting is incorporated into an end-to-end registration network to estimate point saliency scores, enabling the detection of keypoints. Subsequently, the $K$-nearest neighbors method is employed to establish neighborhoods around the keypoints, followed by a permutation-invariant network to extract more detailed descriptors. HRegNet (Lu et al., 2021) is a hierarchical network that leverages geometric features, descriptors, and similarity measures obtained through bilateral consensus and neighborhood consensus to establish correspondence between keypoints. The principles of bilateral consensus and neighborhood consensus suggest that within descriptor space, two correct corresponding points should not only be the nearest neighbors of each other but also exhibit similar neighborhoods. BUFFER (Ao et al., 2023) designs a point-wise learner to enhance computational efficiency and feature representation capabilities by predicting keypoints and estimating point orientations.
@@
Secondly, patches can directly represent the local neighborhood structure. In (Deng et al., 2019), point cloud-FoldNet and point pair features-FoldNet are utilized to extract keypoints from the point cloud patch and obtain permutation-invariant descriptors. Additionally, a new pose estimation method in (Deng et al., 2019) is proposed that achieves faster and more robust results than random sample consensus (RANSAC) (Fischler and Bolles, 1981). StickyPillars (Fischer et al., 2021) integrates keypoint detection and descriptor extraction by jointly learning pixel-level and point-level feature descriptors. YOHO (Wang et al., 2022a) and RoReg (Wang et al., 2023b) employ advanced group equivariant feature learning techniques to achieve rotation invariance, enhancing robustness against variations in point density and noise. Furthermore, the rotation-equivariant component in YOHO and RoReg allows for estimation with just a single correspondence hypothesis, greatly reducing the search space for possible transformations.
@@
Thirdly, voxel grids can achieve uniform sampling of point clouds by adopting grids of different, customizable sizes. 3DSmoothNet (Gojcic et al., 2019) adopts a voxelized smoothed density value technique, incorporating fully convolutional layers to model the local morphology of point clouds. It scrutinizes the local density estimates to accomplish PCR. SpinNet (Ao et al., 2021) eliminates rotational variances by aligning with a reference axis and further reduces them through spherical voxelization and coordinate transformations. It then transforms point clouds into a manageable cylindrical volume and generates representative feature descriptors using cylindrical convolution layers.
@@
Keypoint-free involves considering all potential correspondences rather than detecting critical points. Within this perspective, there exist methods that utilize deep neural networks to directly obtain descriptors that encapsulate vital information. Subsequently, these descriptors are fed into a dedicated module responsible for estimating the transformation parameters. GeDi (Poiesi and Boscaini, 2022) operates by normalizing the local reference frame of the point cloud patch and then encoding it into the descriptors using a deep neural network. These descriptors are invariant to scale and rotation, making them effective for PCR across different application domains.
@@
Other methodologies adopt a coarse-to-refine scheme, in which the matching outcomes are significantly influenced by the descriptors obtained in the initial coarse stage (Qin et al., 2023). GeoTransformer (Qin et al., 2023) encodes the distance and angle information into the transformation representation, enabling effective capture of the geometric structure within individual point clouds and revealing the geometric consistency among the point clouds to be registered. From the viewpoint of considering point-wise and structure, OIF-PCR (Yang et al., 2022) employs an efficient and precise positional encoding strategy during the coarse stage, leveraging a limited number of correspondences. Simultaneously, a joint optimization approach is utilized to optimize the position encoding, progressively refining the point cloud features and reducing the reliance on initialization. RoITr (Yu et al., 2023a) introduces an aggregation module using a rotation invariant Transformer (Vaswani et al., 2017), which is strategically inserted between the encoder and decoder components. Its purpose is to facilitate the extraction of discriminative descriptors that are pose-agnostic and cross-frame position awareness.
@@
3.1.2 Multi-view
The second perspective involves fusing information from multi-view. MVDesc (Zhou et al., 2018) develops a multi-view local descriptor, which is derived from images captured from various viewpoints, specifically for characterizing the 3D keypoints. Subsequently, MVDesc advances a robust matching technique, aimed at rejecting outlier correspondences through efficient belief propagation inference within a defined graphical model. Li et al. (Li et al., 2020b) integrate multi-view rendering into a neural network through a differentiable renderer, allowing the viewpoint to be an optimizable parameter for capturing more informative local context around the interest points. To obtain distinctive descriptors, Li et al. also design a soft view pooling module for fusing convolutional features from different views. Gojcic et al. (Gojcic et al., 2020) utilize iteratively reweighted least squares (IRLS) as a global refinement technique to address the cycle consistency and alleviate the ambiguity of initial alignment in multi-view scanning. However, this approach relies on dense pairwise correspondences, which introduces significant computational overhead and increases the presence of outliers. Consequently, it becomes challenging for IRLS to accurately estimate the correct pose.
To address these limitations, Wang et al. (Wang et al., 2023a) propose a novel approach. They primarily concentrate on learning reliable initialization methods that consider the overlap between multiple point cloud pairs. This enables the construction of sparse yet reliable pose graphs. Furthermore, a history reweighting function is integrated into the IRLS framework, augmenting its generalization and robustness.
@@ 3.2.1 Full-object
The initial category involves full-object PCR, where each point is capable of identifying a unique counterpart in another point cloud. To address this problem, PointNetLK (Aoki et al., 2019) and PointNetLK Revisited (Li et al., 2021), leverage the permutation-invariant network PointNet (Qi et al., 2017) as adaptable imaging functions and integrate them into a recurrent Lucas-Kanade (Lucas and Kanade, 1981) framework. DCP (Wang and Solomon, 2019a) employs graph convolutional neural network and Transformer (Vaswani et al., 2017) modeling to obtain feature representations and capture contextual information. Subsequently, the pointer generation mechanism is employed to estimate the correspondences.
3.2.2 Partial-object
@@
The second category involves partial-object PCR, where not every point has a corresponding point in the other point cloud. Given the common scenario where only a subset of the point clouds to be registered exhibits correspondences, numerous noteworthy studies have emerged in the field of partial-to-partial PCR. These studies have a specific focus on overlap prediction and optimizing the similarity matrix accordingly.
@@
Overlap prediction refers to estimating the overlap region between point clouds to be registered, and then directly finding correspondences in this region. To the best of our knowledge, Predator (Huang et al., 2021a) is the pioneering model that introduces the concept of overlapping region prediction. Predator utilizes a joint encoder and decoder architecture, wherein a graph neural network and an overlap attention module are sequentially applied to enhance contextual relationships and predict the overlap score, respectively. Notably, the overlap attention module facilitates early-stage information interaction in the framework, which positively impacts the estimation of overlapping regions.
@@
With reference to the above concept of information interaction, several approaches are proposed. OMNet (Xu et al., 2021) introduces an innovative mask prediction module that possesses the capability to efficiently generate accurate overlapping masks. Moreover, OMNet establishes a direct connection between the intermediate layers of the mask prediction module and the transformation regression. This connection enables the simultaneous optimization of both the generation of overlapping masks and the estimation of transformation parameters. PCAM (Cao et al., 2021) employs cross-attention matrices (CAM) to achieve feature augmentation. The CAM facilitates simultaneous focus on both shallow geometric information and deep contextual information, enabling the generation of more reliable matching features in overlapping regions.
@@
In addition, several methods enhance the prediction of overlapping regions by employing the Transformer for global modeling. STORM (Wang et al., 2022b) incorporates a differential sampling overlapping prediction module into dual Transformer (Vaswani et al., 2017) layers, which facilitates information exchange between the before and after prediction phases. It employs a dedicated layer that iteratively applies the Gumbel-softmax technique, allowing for the independent sampling of points situated within overlapping regions. REGTR (Yew and Lee, 2022) leverages a main architecture composed of Transformer layers, which incorporate both self-attention and cross-attention mechanisms. These layers are proficient in facilitating the extraction of meaningful and enhanced features. Such an architectural selection empowers the network to accurately predict the probability of each point’s presence in overlapping regions and determine their corresponding positions in another point cloud.
@@
Optimizing the similarity matrix is a crucial aspect of fine-grained correspondence searching. The elements of the similarity matrix indicate the probability of correspondence between individual point pairs. Typically, a probability function is employed to compute the similarity matrix, followed by selecting the maximum value in each row or column to determine the most probable point pairs (Wang and Solomon, 2019a). While softmax is frequently used as the probability function, it tends to produce a blurry correspondence map. To address this issue, numerous studies have emerged to mitigate the ambiguity. PRNet (Wang and Solomon, 2019b) applies the Gumbel-softmax technique to obtain the similarity matrix, a method that finds hard correspondences and alleviates the ambiguity in correspondence search. In addition, to enhance the sharpness of the resultant similarity matrix, a temperature parameter is introduced into the Gumbel-softmax, which can be iteratively adjusted. RPMNet (Yew and Lee, 2020) incorporates the optimal transport layer and annealing to learn a similarity matrix from a hybrid feature composed of spatial coordinates and geometric properties. FIRE-Net (Wu et al., 2021) facilitates feature interactions across various hierarchical levels of point clouds. Initially, FIRE-Net extracts structural features from the point cloud and fosters the interchange of feature information. This process permits points with high feature similarity to effectively perceive each other.

### v2

@@ 1 Introduction
With the progress of sensor technology, acquiring high-precision point cloud data has become more accessible and prevalent Zhang et al. (2023b); Uy et al. (2019). Point cloud registration (PCR), as a pivotal tool in point cloud data processing, aims to align the point cloud data with a common coordinate system, enabling precise three-dimensional (3D) modeling Qin et al. (2023); Liu et al. (2023). This registration process establishes a dependable foundation for point cloud analysis and various applications Huang et al. (2021b).
@@
The differences between our survey and others. Gu et al. (2020) only review traditional PCR methods, DL-based methods are not involved. Zhang et al. (2020) and Huang et al. (2021b) conduct a summary of DL-based PCR methods. However, recent advances in unsupervised methods are not elaborated. Additionally, they did not provide a comprehensive overview of the latest research developments in the PCR field. To address these gaps, we conduct a comprehensive survey and taxonomy of DL-based supervised and unsupervised PCR methods. The taxonomy is summarized in Figure 1, providing a clear and structured overview of the PCR.
@@ 2.1 Definition
PCR Supervised Descriptor Extraction Two-view Multi-view Keypoint-based Keypoint-free Correspondence Search Full-object Partial-object Overlap Prediction Optimizing the Similarity Matrix Outlier Filtering Transformation Parameter Estimation Optimization ICP-based Probabilistic -based Multimodal DeepVCP Lu et al. (2019a), 3DSmoothNet Gojcic et al. (2019), Deng et al. Deng et al. (2019), HRegNet Lu et al. (2021), SpinNet Ao et al. (2021), StickyPillars Fischer et al. (2021), YOHO Wang et al. (2022a), RoReg Wang et al. (2023b), BUFFER Ao et al. (2023). OIF-PCR Yang et al. (2022), GeDi Poiesi and Boscaini (2022), GeoTransformer Qin et al. (2023), RoITr Yu et al. (2023a). MVDesc Zhou et al. (2018), Li et al. Li et al. (2020b), Gojcic et al. Gojcic et al. (2020), Wang et al. Wang et al. (2023a). PointNetLK Aoki et al. (2019), DCP Wang and Solomon (2019a), PointNetLK Revisited Li et al. (2021). Predator Huang et al. (2021a), OMNet Xu et al. (2021), PCAM Cao et al. (2021), STORM Wang et al. (2022b), REGTR Yew and Lee (2022). PRNet Wang and Solomon (2019b), RPMNet Yew and Lee (2020), FIRE-Net Wu et al. (2021) 3DRegNet Pais et al. (2020), DHVR Lee et al. (2021), PointDSC Bai et al. (2021), DLF Zhang et al. (2022a), Chen et al.Chen et al. (2022b), MAC Zhang et al. (2023a). DetarNet Chen et al. (2022c), FINet Xu et al. (2022). DCP Wang and Solomon (2019a), PRNet Wang and Solomon (2019b), IDAM Li et al. (2020a). DeepGMR Yuan et al. (2020), OGMM Mei et al. (2023a), Chen et al.Chen et al. (2023a), VBReg Jiang et al. (2023). PCR-CG Zhang et al. (2022c), ImLoveNet Chen et al. (2022a), IMFNET Huang et al. (2022c), GMF Huang et al. (2022b), PEAL Yu et al. (2023b). Unsupervised Correspondence -free PPF-FoldNet Deng et al. (2018), PCRNet Sarode et al. (2019), UPCR Zhang et al. (2021), UGMM Huang et al. (2022a), Sun et al. Sun et al. (2023). Correspondence -based CEMNet Jiang et al. (2021), RIENet Shen et al. (2022), UDPReg Mei et al. (2023b).
Figure 1: A taxonomy of PCR algorithms.
@@ 2.2 Datasets
Datasets for PCR can be broadly classified into two categories: artificially synthesized and acquired through real instruments. Each category exhibits unique characteristics and has different levels of applicability in PCR tasks. Synthesized point cloud datasets are typically composed of virtual models created by using computer graphics techniques to replicate real-world environments. These datasets contain the object-level ModelNet40 Wu et al. (2015) and ShapeNet Chang et al. (2015), which consist of data generated through computer-aided design, as well as the scene-level dataset ICL-NUIM Choi et al. (2015) and FlyingShapes Chen et al. (2023b).
However, while synthetic data are beneficial for certain applications, they often lack the complexity and variability found in real-world scenarios. Consequently, incorporating real-world data into training and validation processes is essential for obtaining robust algorithms. Datasets comprising realistic point cloud data include Stanford Curless and Levoy (1996), ETH Pomerleau et al. (2012), KITTI Geiger et al. (2012), Apollo-SouthBay Lu et al. (2019b), ScanObjectNN Uy et al. (2019), and WHU-TLS Dong et al. (2020). Furthermore, there is a 3DMatch Zeng et al. (2017) dataset that comprises both synthesized and realistic scans. The attributes of various datasets are summarized in Table 1.
@@ 2.3 Metrics
Stanford Curless and Levoy (1996) Real 10 O
ETH Pomerleau et al. (2012) Real 36 S
KITTI Geiger et al. (2012) Real 22 S
ModelNet40 Wu et al. (2015) Syn 12311 O
ShapeNet Chang et al. (2015) Syn 55000+ O
ICL-NUIM Choi et al. (2015) Syn 8 S
3DMatch Zeng et al. (2017) Syn&Real 62 S
Apollo-SouthBay Lu et al. (2019b) Real 6 S
ScanObjectNN Uy et al. (2019) Real 15000+ O
WHU-TLS Dong et al. (2020) Real 115 S
FlyingShapes Chen et al. (2023b) Syn 200 S
Table 1: Datasets for PCR task. Syn means synthetic point clouds. Real means realistic point clouds. S and O denote scene-level and object-level, respectively.
@@ 3.1 Descriptor Extraction
Two-view. The first perspective involves two-view registration, which emerges as the prevalent approach in the field of PCR. We further classify these methods into two categories: keypoint-based and keypoint-free.
@@
Firstly, points, serving as the fundamental elements of point clouds, are discrete and unlinked 3D entities. Consequently, extracting descriptors from points typically necessitates the construction of intricate local relationships. In DeepVCP Lu et al. (2019a), point weighting is incorporated into an end-to-end registration network to estimate point saliency scores, enabling the detection of keypoints. Subsequently, the $K$-nearest neighbors method is employed to establish neighborhoods around the keypoints, followed by a permutation-invariant network to extract more detailed descriptors. HRegNet Lu et al. (2021) is a hierarchical network that leverages geometric features, descriptors, and similarity measures obtained through bilateral consensus and neighborhood consensus to establish correspondence between keypoints. The principles of bilateral consensus and neighborhood consensus suggest that within descriptor space, two correct corresponding points should not only be the nearest neighbors of each other but also exhibit similar neighborhoods. BUFFER Ao et al. (2023) designs a point-wise learner to enhance computational efficiency and feature representation capabilities by predicting keypoints and estimating point orientations.
@@
Secondly, patches can directly represent the local neighborhood structure. In Deng et al. (2019), point cloud-FoldNet and point pair features-FoldNet are utilized to extract keypoints from the point cloud patch and obtain permutation-invariant descriptors. Additionally, a new pose estimation method in Deng et al. (2019) is proposed that achieves faster and more robust results than random sample consensus (RANSAC) Fischler and Bolles (1981). StickyPillars Fischer et al. (2021) integrates keypoint detection and descriptor extraction by jointly learning pixel-level and point-level feature descriptors. YOHO Wang et al. (2022a) and RoReg Wang et al. (2023b) employ advanced group equivariant feature learning techniques to achieve rotation invariance, enhancing robustness against variations in point density and noise. Furthermore, the rotation-equivariant component in YOHO and RoReg allows for estimation with just a single correspondence hypothesis, greatly reducing the search space for possible transformations.
@@
Thirdly, voxel grids can achieve uniform sampling of point clouds by adopting grids of different, customizable sizes. 3DSmoothNet Gojcic et al. (2019) adopts a voxelized smoothed density value technique, incorporating fully convolutional layers to model the local morphology of point clouds. It scrutinizes the local density estimates to accomplish PCR. SpinNet Ao et al. (2021) eliminates rotational variances by aligning with a reference axis and further reduces them through spherical voxelization and coordinate transformations. It then transforms point clouds into a manageable cylindrical volume and generates representative feature descriptors using cylindrical convolution layers.
@@
Keypoint-free involves considering all potential correspondences rather than detecting critical points. Within this perspective, there exist methods that utilize deep neural networks to directly obtain descriptors that encapsulate vital information. Subsequently, these descriptors are fed into a dedicated module responsible for estimating the transformation parameters. GeDi Poiesi and Boscaini (2022) operates by normalizing the local reference frame of the point cloud patch and then encoding it into the descriptors using a deep neural network. These descriptors are invariant to scale and rotation, making them effective for PCR across different application domains.
@@
Other methodologies adopt a coarse-to-refine scheme, in which the matching outcomes are significantly influenced by the descriptors obtained in the initial coarse stage Qin et al. (2023). GeoTransformer Qin et al. (2023) encodes the distance and angle information into the transformation representation, enabling effective capture of the geometric structure within individual point clouds and revealing the geometric consistency among the point clouds to be registered. From the viewpoint of considering point-wise and structure, OIF-PCR Yang et al. (2022) employs an efficient and precise positional encoding strategy during the coarse stage, leveraging a limited number of correspondences. Simultaneously, a joint optimization approach is utilized to optimize the position encoding, progressively refining the point cloud features and reducing the reliance on initialization. RoITr Yu et al. (2023a) introduces an aggregation module using a rotation invariant Transformer Vaswani et al. (2017), which is strategically inserted between the encoder and decoder components. Its purpose is to facilitate the extraction of discriminative descriptors that are pose-agnostic and cross-frame position awareness.
@@
Multi-view. The second perspective involves fusing information from multi-view. MVDesc Zhou et al. (2018) develops a multi-view local descriptor, which is derived from images captured from various viewpoints, specifically for characterizing the 3D keypoints. Subsequently, MVDesc advances a robust matching technique, aimed at rejecting outlier correspondences through efficient belief propagation inference within a defined graphical model. Li et al. Li et al. (2020b) integrate multi-view rendering into a neural network through a differentiable renderer, allowing the viewpoint to be an optimizable parameter for capturing more informative local context around the interest points. To obtain distinctive descriptors, Li et al. also design a soft view pooling module for fusing convolutional features from different views. Gojcic et al. Gojcic et al. (2020) utilize iteratively reweighted least squares (IRLS) as a global refinement technique to address the cycle consistency and alleviate the ambiguity of initial alignment in multi-view scanning. However, this approach relies on dense pairwise correspondences, which introduces significant computational overhead and increases the presence of outliers. Consequently, it becomes challenging for IRLS to accurately estimate the correct pose.
To address these limitations, Wang et al. Wang et al. (2023a) propose a novel approach. They primarily concentrate on learning reliable initialization methods that consider the overlap between multiple point cloud pairs. This enables the construction of sparse yet reliable pose graphs. Furthermore, a history reweighting function is integrated into the IRLS framework, augmenting its generalization and robustness.
@@ 3.2 Correspondence Search
Full-object. The initial category involves full-object PCR, where each point is capable of identifying a unique counterpart in another point cloud. To address this problem, PointNetLK Aoki et al. (2019) and PointNetLK Revisited Li et al. (2021), leverage the permutation-invariant network PointNet Qi et al. (2017) as adaptable imaging functions and integrate them into a recurrent Lucas-Kanade Lucas and Kanade (1981) framework. DCP Wang and Solomon (2019a) employs graph convolutional neural network and Transformer Vaswani et al. (2017) modeling to obtain feature representations and capture contextual information. Subsequently, the pointer generation mechanism is employed to estimate the correspondences.
@@
Partial-object. The second category involves partial-object PCR, where not every point has a corresponding point in the other point cloud. Given the common scenario where only a subset of the point clouds to be registered exhibits correspondences, numerous noteworthy studies have emerged in the field of partial-to-partial PCR. These studies have a specific focus on overlap prediction and optimizing the similarity matrix accordingly.
@@
Overlap prediction refers to estimating the overlap region between point clouds to be registered, and then directly finding correspondences in this region. To the best of our knowledge, Predator Huang et al. (2021a) is the pioneering model that introduces the concept of overlapping region prediction. Predator utilizes a joint encoder and decoder architecture, wherein a graph neural network and an overlap attention module are sequentially applied to enhance contextual relationships and predict the overlap score, respectively. Notably, the overlap attention module facilitates early-stage information interaction in the framework, which positively impacts the estimation of overlapping regions.
@@
With reference to the above concept of information interaction, several approaches are proposed. OMNet Xu et al. (2021) introduces an innovative mask prediction module that possesses the capability to efficiently generate accurate overlapping masks. Moreover, OMNet establishes a direct connection between the intermediate layers of the mask prediction module and the transformation regression. This connection enables the simultaneous optimization of both the generation of overlapping masks and the estimation of transformation parameters. PCAM Cao et al. (2021) employs cross-attention matrices (CAM) to achieve feature augmentation. The CAM facilitates simultaneous focus on both shallow geometric information and deep contextual information, enabling the generation of more reliable matching features in overlapping regions.
@@
In addition, several methods enhance the prediction of overlapping regions by employing the Transformer for global modeling. STORM Wang et al. (2022b) incorporates a differential sampling overlapping prediction module into dual Transformer Vaswani et al. (2017) layers, which facilitates information exchange between the before and after prediction phases. It employs a dedicated layer that iteratively applies the Gumbel-softmax technique, allowing for the independent sampling of points situated within overlapping regions. REGTR Yew and Lee (2022) leverages a main architecture composed of Transformer layers, which incorporate both self-attention and cross-attention mechanisms. These layers are proficient in facilitating the extraction of meaningful and enhanced features. Such an architectural selection empowers the network to accurately predict the probability of each point’s presence in overlapping regions and determine their corresponding positions in another point cloud.
@@
Optimizing the similarity matrix is a crucial aspect of fine-grained correspondence searching. The elements of the similarity matrix indicate the probability of correspondence between individual point pairs. Typically, a probability function is employed to compute the similarity matrix, followed by selecting the maximum value in each row or column to determine the most probable point pairs Wang and Solomon (2019a). While softmax is frequently used as the probability function, it tends to produce a blurry correspondence map. To address this issue, numerous studies have emerged to mitigate the ambiguity. PRNet Wang and Solomon (2019b) applies the Gumbel-softmax technique to obtain the similarity matrix, a method that finds hard correspondences and alleviates the ambiguity in correspondence search. In addition, to enhance the sharpness of the resultant similarity matrix, a temperature parameter is introduced into the Gumbel-softmax, which can be iteratively adjusted. RPMNet Yew and Lee (2020) incorporates the optimal transport layer and annealing to learn a similarity matrix from a hybrid feature composed of spatial coordinates and geometric properties. FIRE-Net Wu et al. (2021) facilitates feature interactions across various hierarchical levels of point clouds. Initially, FIRE-Net extracts structural features from the point cloud and fosters the interchange of feature information. This process permits points with high feature similarity to effectively perceive each other.